Circular No. 9112

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V5585 SAGITTARII = NOVA SAGITTARII 2010 John Seach, Chatsworth Island, NSW, Australia, reports his discov-

ery of an apparent nova (mag 8.5) on three 6-s CCD exposures (limiting mag 11) obtained on Jan. 20.72 UT (time corrected from that published on CBET 2140) with a Nikon D300 camera (+ 50-mm f/1.4 lens + orange filter), providing a position for the variable as α = 18^{h}07^{m}26^{s}, δ = −29^{o}01′ (equinox 2000.0); nothing is visible at this position on his images taken on 2009 Nov. 15.89 (limiting mag 11.5). Following posting on the Central Bureau’s unconfirmed-objects webpage, S. Kiyota (Tsukuba, Japan) reports that CCD images taken on Jan. 22 with a 20-cm reflector (and Apogee Alta E47+ CCD camera) yield the following position end figures for the variable: 26^{s}.79, 00′42′′.8. Kiyota also provides the following magnitudes for the nova: Jan. 22.862, V = 8.98, R_{c} = 8.38; 22.864, I_{c} = 7.78. W. Vollmann, Vienna, Austria, reports that he measured green mag 9.3 from unfiltered images taken by H. Koberger (Vilaflor, Tenerife) with a digital SLR camera (+ 300-mm camera lens) on Jan. 22.287, adding that nothing is visible within 10′′ of its position on Digitized Sky Survey blue and red images to limiting mag ≈ 18. K. Nishiyama (Kurume, Fukuoka-ken, Japan) and F. Kabashima (Miyaki-cho, Saga-ken, Japan) report mag 8.5 and position end figures 26^{s}.95, 00′43′′.6 from five 5-s unfiltered CCD frames (limiting mag 15.0) taken around Jan. 23.892 using a Meade 0.40-m reflector (+ SBIG STL1001E camera). H. Maehara, Kwasan Observatory, Kyoto University, reports that a low-resolution spectrogram (range 400–800 nm), obtained on Jan. 23.887 with the 25-cm telescope of the Kwasan Observatory, shows Hα, Hβ, and Fe II emission lines, suggesting that this object is a “Fe II”-class nova. Additional details are given on CBETs 2140 and 2142. E. Kazarovets and N. N. Samus report that the GCVS designation V5585 Sgr has been assigned to this nova.

COMETS C/2009 M4 AND C/2009 M5 (SOHO) Further to IAUC 9094, additional Kreutz sungrazers have been found

on SOHO website images. C/2009 M4 was small and stellar in appearance (peak mag ∼ 7) on C3 images, but quite diffuse (mag ∼ 8.5) and elongated in C2 images. C/2009 M5 was very faint (mag ∼ 8.5) and very diffuse.

Comet 2009 UT α_{2000} δ_{2000} Inst. F MPEC
C/2009 M4 June 24.488 6^{h}04.^{m}5 +21^{o}06′ C3/2 RK 2009-P02
C/2009 M5  25.038 6 07.6 +21 50 C2  BZ 2009-P02
2010 January 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green